OpenStudy (loser66):

How to prove: \(||x+y||||x-y||\leq ||x||^2 +||y||^2\) Please, help

OpenStudy (loser66):

Triangle inequality shows: \[||x+y||\leq||x||+||y||\] But I don't know the case ||x -y||

OpenStudy (loser66):

mood up, right? hihihi

OpenStudy (ikram002p):

|x-y| |x+y|=|x^2-y^2| this should work

OpenStudy (ikram002p):

|x^2-y^2| <=|x^2+y^2|<=|x^2|+|y^2|

OpenStudy (loser66):

It works, but we have to use inequality theorem to show, friend, or Cauchy Schwarz

OpenStudy (ikram002p):

triangle is enough i guess

OpenStudy (loser66):

ok, it works,:) thank you, friend

OpenStudy (ikram002p):

wait , but what if u prof want it as what u said ? u cant close the question :O

OpenStudy (loser66):

why not? it makes sense to me,now I can use it as you said, clearly, |x|^2 -|y|^2 < |x|^2+|y|^2

OpenStudy (loser66):

and the left hand side satisfies the triangle inequality, We just make them link by skipping the middle term.

OpenStudy (ikram002p):

yep ^^